A nearsighted man uses his ideally prescribed eyeglasses that have a refractive power of −3.50 diopters. He would like a prescription for contact lenses. What is the correct contact lens prescription if he normally wears his eyeglasses a distance of 2.2 cm from his eyes?
Answer #1

Lets find near point of his eyes,

-1/[d-2.2] = 1/f

d = -f+2.2

= 100/3.5 + 2.2

= 30.77 cm

for contact lens, object distance will be 30.77 cm = 0.3077 m

Power = -1/0.3077 = -3.25 D